OpenStudy (anonymous):

Write the equation y = 2 - 4x in standard form. A.y - 2 = 4(x - 0) B.4x - y = -2 C.4x + y = 2 D.4x + 2y = 0

OpenStudy (anonymous):

Well, do you know what Standard Form is?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

Standard Form is: \[Ax + By = C\]

OpenStudy (anonymous):

Now, to get the 4x in the front, you just have to add 4x to both sides.

OpenStudy (anonymous):

Do you know how to solve it from there?

OpenStudy (anonymous):

ok. I still don't quit get how to solve it... :/

OpenStudy (anonymous):

You have y = 2 - 4x Now, you add 4x to both sides: y = 2 - 4x +4x + 4x Now what does the final equation look like?

OpenStudy (anonymous):

Right right. so it would be 4x+y=2 @wmckinely

OpenStudy (anonymous):

Yup:) Good job! :D